### v1

# Optimal Control on Positive Cones

Abstract
An optimal control problem on finite-dimensional positive cones is stated. Under a critical assumption on the cone, the corresponding Bellman equation is satisfied by a linear function, which can be computed by convex optimization. A separate theorem relates the assumption on the cone to the existence of minimal elements in certain subsets of the dual cone. Three special cases are derived as examples. The first one, where the positive cone is the set of positive semi-definite matrices, reduces to standard linear quadratic control. The second one, where the positive cone is a polyhedron, reduces to a recent result on optimal control of positive systems. The third special case corresponds to linear quadratic control with additional structure, such as spatial invariance.
1 Introduction
The theory of optimal control has been a cornerstone of control theory ever since the early work by Bellman [1]. Under standard assumptions, a discrete time optimal control problem with non-negative cost
Minimize $\sum_{t=0}^{\infty}g(x(t),u(t))$
subject to $x(t+1)=f(x(t),u(t)),\quad x(0)=x_{0}$
$u(t)\in U(x(t))$
is solvable if and only if the Bellman equation
$J^{*}(x)$ $=\min_{u\in U(x)}\left[g(x,u)+J^{*}(f(x,u))\right]$
has a non-negative solution $J^{*}$ which is finite for all $x$. The minimizing $u$ gives an optimal control policy. See for example [2, section 3.1].
Of particular importance are the rare situations when the Bellman equation has an explicit solution. The most well known case is probably linear dynamics and quadratic cost, introduced in the pioneering work of [3]. Another case of fundamental importance is when both $x$ and $u$ are restricted to finite sets. This setting leads to shortest path problems on graphs, with an extensive literature dating back to the work of [4]. Recently, a third class of problems with explicit solutions to the Bellman equation was introduced in the context of linear systems on the positive orthant in $\mathbb{R}^{n}$, i.e. so called positive systems, see [5, 6, 7].
The purpose of this note is to show that the previous instances where the Bellman equation can be solved explicitly are in fact special cases of a more general class of optimal control problems stated in terms of positive cones.
2 Main Result
Theorem 1.
Let $\mathcal{P}$ be a positive cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$. Given $(s,r)\in\mathcal{P}^{*}$ and $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, suppose that for every $(x,u)\in\mathcal{P}$ there exists $v$ such that $(Ax+Bu,v)\in\mathcal{P}$. Moreover, assume the existence of $\phi:\mathbb{R}^{m}\to\mathbb{R}^{n}$ such that
$\phi(\mu)^{\top}x$ $=\min_{(x,u)\in\mathcal{P}}\mu^{\top}u$ (1)
for all $(x,v)\in\mathcal{P}$, $(\lambda,\mu)\in\mathcal{P}^{*}$. Then the following three statements are equivalent:
  • ($i$)
    The following optimal control problem has a finite value for every $(x_{0},u_{0})\in\mathcal{P}$:
    Minimize $\sum_{t=0}^{\infty}[s^{\top}x(t)+r^{\top}u(t)]\text{ over }\{u(t)\}_{t=0}^{\infty}$
    subject to $x(t+1)=Ax(t)+Bu(t)$
    $(x(t),u(t))\in\mathcal{P},\quad x(0)=x_{0}.$
  • ($ii$)
    There exists $(\lambda_{*},0)\in\mathcal{P}^{*}$ such that
    $\lambda_{*}=s+A^{\top}\lambda_{*}+\phi(r+B^{\top}\lambda_{*}).$ (2)
  • ($iii$)
    The convex program
    Maximize $\lambda^{\top}x_{0}\text{ over }(\lambda,0)\in\mathcal{P}^{*}$
    subject to $(s+A^{\top}\lambda-\lambda,r+B^{\top}\lambda)\in\mathcal{P}^{*}.$
    has a bounded value for every $(x_{0},u_{0})\in\mathcal{P}$.
If ($ii$) is true, then the minimal value of ($i$) and the maximal value of ($iii$) are both equal to $\lambda_{*}^{\top}x_{0}$. Moreover, the optimal control law $x\mapsto u$ is obtained by minimizing $(r+B^{\top}\lambda_{*})^{\top}u$ over $u$ subject to the constraint $(x,u)\in\mathcal{P}$.
Proof.
The minimization in ($i$) is a special case of the general formulation in the introduction, where $f(x,u)=Ax+Bu$ and $g(x,u)=s^{\top}x+r^{\top}u$. Define the sequence $\{\lambda_{k}\}_{k=0}^{\infty}$ recursively by
$\lambda_{k+1}$ $:=s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k}),$ $\lambda_{0}$ $:=0.$
If $J_{k}(x)=\lambda_{k}^{\top}x$, then
$\min_{u\in U(x)}\left[g(x,u)+J_{k}(f(x,u))\right]$
$=\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)\right]$
$=s^{\top}x+\lambda_{k}^{\top}Ax+\min_{(x,u)\in\mathcal{P}}\left[(r+B^{\top}\lambda_{k})^{\top}u\right]$
$=\left[s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k})\right]^{\top}x$
$=\lambda_{k+1}^{\top}x.$
Hence, value iteration
$J_{k+1}(x)$ $=\min_{u\in U(x)}\left[g(x,u)+J_{k}(f(x,u))\right],$ $J_{0}$ $=0$
gives a sequence $0=J_{0}\leq J_{1}\leq J_{2}\leq\ldots$ with $J_{k}(x)=\lambda_{k}^{\top}x$ for all $k$. In particular $(\lambda_{k},0)\in\mathcal{P}_{*}$ for all $k$. If ($i$) holds, then the sequence has an upper limit $J^{*}(x)=\lambda_{*}^{\top}x$ with $\lambda_{*}$ satisfying the condition in ($ii$). Conversely, if ($ii$) holds, then $(\lambda_{*}-\lambda_{k},0)\in\mathcal{P}_{*}$ for all $k$, so the sequence $J_{0}\leq J_{1}\leq J_{2}\leq\ldots$ has an upper limit and ($i$) holds too. This proves the equivalence of ($i$) and ($ii$), as well as the formula $\lambda_{*}^{\top}x_{0}$ for the minimal value in ($i$). The optimal controller is $u(t)=\lambda(x(t)$, where
$\lambda(x)$ $=\arg\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{*}^{\top}(Ax+Bu)\right]$
$=\arg\min_{(x,u)\in\mathcal{P}}\left[(r+B^{\top}\lambda_{*})^{\top}u\right].$
It remains to prove equivalence betwen ($ii$) and ($iii$).
First, by definition of $\phi$, the optimization constraint in ($iii$) implies that
$\lambda^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda^{\top}(Ax+Bu)$ (3)
for all $(x,u)\in\mathcal{P}$. It is a well known fact that solutions to the Bellman inequality are lower bounds on the optimal cost. Hence, if ($ii$) holds, the maximum for the convex program must be attained at the solution of eq. 2, so also ($iii$) holds. Conversely, suppose that ($iii$) holds with the maximizing argument $\lambda_{k}$. Define
$\lambda_{k+1}$ $:=s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k}).$
It follows from eq. 3 that
$\lambda_{k}^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)$
$\lambda_{k}^{\top}x$ $\leq s^{\top}x+\lambda_{k}^{\top}Ax+\phi(r+B^{\top}\lambda_{k})^{\top}x$
for all $(x,u)\in\mathcal{P}$ and therefore
$(\lambda_{k+1}-\lambda_{k},0)$ $=(s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k})-\lambda_{k},0)\in\mathcal{P}_{*}.$
It follows that
$\lambda_{k+1}^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda_{k+1}^{\top}(Ax+Bu)$
Hence also $\lambda_{k+1}$ satisfies the constraints of the optimization in ($iii$), so $\lambda_{k+1}=\lambda_{k}$. Otherwise $\lambda_{k}$ would not be maximizing. This means that $\lambda_{k}$ solves eq. 2 and also ($ii$) holds. The proof is complete. ∎
1 supposes the existence of a function $\phi$ with the key property that it renders the optimization in eq. 1 linear in $x$. It is this feature that keeps the value iteration linear in $x$, which is instrumental in allowing the recasting of the optimal control problem as either a fixed point equation or convex program (conditions ($ii$) and ($iii$) in 1). It is natural then to ask when such a $\phi$ exists. The following result shows that when the cone $\mathcal{P}$ is proper, this can be established in a rather direct way through a property of the dual cone $\mathcal{P}^{*}$. We will use this result in the next section when we connect 1 to some well studied optimal control problems.
Theorem 2.
Let $\mathcal{P}$ be a proper cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$, $\lambda^{*}\in\mathbb{R}^{n}$, and $\mu\in\mathbb{R}^{m}$. Suppose that there exists a $\lambda\in\mathbb{R}^{n}$ such that
$\left(\lambda,\mu\right)\in\operatorname{rel}\operatorname{int}\mathcal{P}^{*}.$ (4)
Then the following statements are equivalent:
  1. ($i$)
    $-\lambda^{*\top{}}x=\inf_{\left(x,u\right)\in\mathcal{P}}\mu^{\top}u.$ (5)
  2. ($ii$)
    $\left(\lambda^{*},\mu\right)\in\mathcal{P}^{*},\;\text{and for all}\;\left(\lambda,\mu\right)\in\mathcal{P}^{*},\;\left(\lambda-\lambda^{*},0\right)\in\mathcal{P}^{*}.$
Proof.
($i$$\implies$($ii$): With a view to generating a contradiction, suppose that under the hypothesis of ($i$),
$\left(\lambda^{*},\mu\right)\notin\mathcal{P}^{*}.$
There therefore exists an
$\left(\bar{x},\bar{u}\right)\in\mathcal{P}\;\text{such that}\;\begin{bmatrix}\lambda^{*}\\ \mu\end{bmatrix}^{\top}\begin{bmatrix}\bar{x}\\ \bar{u}\end{bmatrix}<0,$
which implies that $-\lambda^{*\top{}}\bar{x}>\mu^{\top}\bar{u}$. However from eq. 5 we see that $-\lambda^{*\top{}}\bar{x}\leq{}\mu^{\top}\bar{u}$, which is a contradiction. This proves the first claim in ($ii$). Now, again with a view to generating a contradiction, suppose that
$\left(\lambda,\mu\right)\in\mathcal{P}^{*}\;\text{and}\;\left(\lambda-\lambda^{*},0\right)\notin\mathcal{P}^{*}.$
Therefore there exists an
$\left(\tilde{x},\tilde{u}\right)\in\mathcal{P}\;\text{such that}\;\begin{bmatrix}\lambda-\lambda^{*}\\ 0\end{bmatrix}^{\top}\begin{bmatrix}\tilde{x}\\ \tilde{u}\end{bmatrix}<0,$
which implies that $\lambda^{*\top{}}\tilde{x}>\lambda^{\top}\tilde{x}$. The dual of eq. 5 reads
$\sup_{\left(\lambda,\mu\right)\in\mathcal{P}^{*}}-\lambda^{\top}x.$ (6)
The solution to eq. 6 is lower bounded by $-\lambda^{\top}\tilde{x}$. However by weak duality, the solution to eq. 6 is upper bounded by $-\lambda^{*\top{}}\tilde{x}$. Therefore
$\left(\lambda,\mu\right)\in\mathcal{P}^{*}\implies\lambda^{*\top{}}\tilde{x}\leq{}\lambda^{\top}\tilde{x}.$
Again this a contradiction, proving the claim.
($ii$$\implies$($i$): Since $\mathcal{P}$ is a proper cone, $\mathcal{P}^{**}=\mathcal{P}$. Therefore for all $\left(x,u\right)\in\mathcal{P}$,
$\begin{bmatrix}\lambda-\lambda^{*}\\ 0\end{bmatrix}^{\top}\begin{bmatrix}x\\ u\end{bmatrix}\geq{}0,\implies-\left(\lambda-\lambda^{*}\right)^{\top}x\leq{}0.$
Therefore for any such $\lambda$,
$-\lambda^{\top}x=-\left(\lambda-\lambda^{*}\right)^{\top}x-\lambda^{*\top{}}x\leq{}-\lambda^{*\top{}}x.$
It then follows that the solution to eq. 6 equals $-\lambda^{*\top{}}x$. Condition eq. 4 states that eq. 6 is strictly feasible. This implies that strong duality holds, which proves eq. 5 as required. ∎
3 Examples
3.1 The Linear Quadratic Regulator
In this example we will connect the standard linear-quadratic-regulator (LQR) problem to 1. This will give an interpretation of our more abstract notation and results in a hopefully more familiar setting. In particular we will show that when the optimization in ($i$) corresponds to the LQR problem, conditions ($ii$) and ($iii$) reduce to the algebraic Riccati equations and semi-definite programs that are standard in the solution to this problem.
We will in fact study a slightly more general problem. The connections to the standard LQR will be established at the end of the section. The optimization problem that we consider first is:
Minimize $\sum_{t=0}^{\infty}\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)$
over $\{U_{1}(t),U_{2}(t)\}_{t=0}^{\infty},$
subject to $X(t+1)=\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix},$
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\succeq{}0,\;X(0)=X_{0}\succeq{}0.$
In the above, the variables are the matrices $X(t)$, $U_{1}(t)$ and $U_{2}(t)$. The problem data is the matrices $S$, $R_{1}$, $R_{2}$, $F$ and $G$, where we additionally require that
$\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\succeq{}0.$ (7)
Although it may not be immediately clear at first sight, this is an optimization of precisely the type considered in 1($i$) where $\mathcal{P}$ is the positive semi-definite cone (which is positive, proper and self-dual).
We will now explain the correspondence between the variables and data in the above, and those in $1$The set of $p\times{}p$ positive semi-definite matrices is a proper cone in $\mathbb{R}^{p\left(p+1\right)/2}$, and so the mappings required to match the two problems explicitly can certainly be defined. However manipulations on the semi-definite cone are often more conveniently expressed in terms of matrices. We will have a need for such manipulations to identify the function $\phi$, and so proceed with this more informal connection between the two problems.. The variables are related as follows:
$x(t)$ $\longleftrightarrow{}X(t)$
$u(t)$ $\longleftrightarrow{}\begin{bmatrix}0&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}(t)\end{bmatrix}.$
That is the $x(t)$ and $u(t)$ variables correspond to a partitioning of a larger matrix. For the dynamics we have that:
$Ax(t)$ $\longleftrightarrow{}FX(t)F^{\top}$
$Bu(t)$ $\longleftrightarrow{}\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}0&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix}.$
These dyanamics are clearly cone preserving under the constraint $\left(x,u\right)\in\mathcal{P}$, as required by 1. Similarly the costs can be connected through:
$s^{\top}x(t)$ $\longleftrightarrow{}\operatorname{tr}\left(SX(t)\right)$
$r^{\top}u(t)$ $\longleftrightarrow{}\operatorname{tr}\left(\begin{bmatrix}0&R_{1}\\ R_{1}^{\top}&R_{2}\end{bmatrix}\begin{bmatrix}0&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right).$
Since the semi-definite cone is self-dual, under eq. 7 we have that $\left(s,r\right)\in\mathcal{P}^{*}$ as required. We have now established that our matrix optimization problem is on the form of 1($i$), and that all the prerequisits of the Theorem are satisfied except for the existence of the function $\phi$. We will now address this using 2.
To this end, let us now examine 2 when $\mathcal{P}$ is the semi-definite cone. Again it is helpful to relate vector and matrix variables, and so introduce
$\lambda$ $\longleftrightarrow{}\Lambda$
$\mu$ $\longleftrightarrow{}\begin{bmatrix}0&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}.$
After specialising the notation to this setting, 2 then states that the existence of $\phi\left(\mu\right)$ is equivalent to the set
$\mathcal{C}_{\mu}=\left\{\Lambda:\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\succeq{}0\right\},$
having a minimum element (that is, an element $\Lambda^{*}\in\mathcal{C}_{\mu}$ such that $\Lambda^{*}\preceq{}\Lambda$ for all $\Lambda\in\mathcal{C}_{\mu}$). For the semi-definite cone whenever $\mathcal{C}_{\mu}$ is non-empty, this element exists, and is given by $\Lambda^{*}=M_{1}M_{2}^{+}M_{1}^{\top}{}$. Therefore
$\phi\left(\mu\right)\longleftrightarrow{}-M_{1}M_{2}^{+}M_{1}^{\top}{}.$
We are now ready to examine 1 ($ii$)–($iii$). Substituting in for all found correspondances, we see that in our matrix notation [^2] It is perhaps worth noting that from basic properties of adjoints, it follows that $A^{\top}\lambda(t)$ $\longleftrightarrow{}F^{\top}\Lambda(t)F$ $B^{\top}\lambda(t)$ $\longleftrightarrow{}\begin{bmatrix}0&F^{\top}\Lambda(t)G\\ G^{\top}\Lambda(t)F&G^{\top}\Lambda(t)G\end{bmatrix}.$ , the equation $\lambda_{*}=s+A^{\top}\lambda_{*}+\phi(r+B^{\top}\lambda_{*})$ becomes
$\Lambda^{*}=S+F^{\top}\Lambda^{*}F-\left(F^{\top}\Lambda^{*}G+R_{1}\right)\left(R_{2}+G^{\top}{}\Lambda^{*}G\right)^{+}\left(G^{\top}\Lambda^{*}F+R_{1}^{\top}{}\right).$
Note that a degree of care has to be taken when reading this formula, since the requirements on the domain of $\phi$ have been left implicit. For example, these are guaranteed to hold under the common assumption that $R_{2}\succ{}0$, after which the above collapses to the usual algebraic Riccati equation. This can then be solved in the usual way whenever the pair $\left(F,G\right)$ is stabilizable. 1 ($iii$) can be similarly analyzed. In this case, the convex program reduces to the semi-definite program
Maximize $\operatorname{tr}\left(\Lambda{}X_{0}\right)$
subject to $\begin{bmatrix}S-\Lambda&R_{1}\\ R_{1}^{\top}&R_{2}\end{bmatrix}+\begin{bmatrix}F^{\top}{}\\ G^{\top}{}\end{bmatrix}\Lambda\begin{bmatrix}F&G\end{bmatrix}\succeq{}0,\;\Lambda\succeq{}0.$
To conclude the example, let us finally connect explicitly to the standard LQR problem. The standard LQR problem is concerned with the optimization
Minimize $\sum_{t=0}^{\infty}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}^{\top}\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}\text{ over }\{u(t)\}_{t=0}^{\infty}$
subject to $x(t+1)=Fx(t)+Gu(t),\,x(0)=x_{0}.$
To see the connection to the matrix optimization only requires us to notice that whenever
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}=\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}\begin{bmatrix}x(t)^{\top}&u(t)^{\top}\end{bmatrix},$ (8)
the specified matrix dynamics ensure that
$X(t+1)=x(t+1)x(t+1)^{\top}{},$
and also that
$\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)=$
$\qquad\qquad\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}^{\top}\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}.$
That is, under eq. 8, the costs and constraints in the standard LQR problem, and the matrix optimization problem, are identical. It then follows (with a little extra work) that the standard LQR problem corresponds to the special case of the matrix optimization when the initial condition $X_{0}$ is rank 1.
3.2 A Linear Regulator on a Polyhedral Cone
In this subsection we consider the application of 1 when $\mathcal{P}$ is the polyhedral cone (which are always positive and proper)
$\mathcal{P}=\left\{\begin{bmatrix}x\\ u\end{bmatrix}:\begin{bmatrix}I&0\\ E&-I\\ E&I\end{bmatrix}\begin{bmatrix}x\\ u\end{bmatrix}\geq{}0\right\},$
where $E\geq{}0$. The motivation for this choice comes from the study of positive systems. Note in particular that the condition $(x,u)\in\mathcal{P}$ constrains the system state to lie in the positive orthant, and the input to satisfy
$\left|{u}\right|\leq{}Ex.$
When minimising the cost in 1 ($i$), we are in effect optimizing performance under the constraint that the control keeps the system state positive. In what follows will recover the results of [5, Theorem 1].
First note that the dual cone $\mathcal{P}^{*}$ is also polyhedral, and is given by
$\mathcal{P}^{*}=\left\{\begin{bmatrix}\lambda\\ \mu\end{bmatrix}:\begin{bmatrix}\lambda\\ \mu\end{bmatrix}=\begin{bmatrix}I&0\\ E&-I\\ E&I\end{bmatrix}^{\top}\begin{bmatrix}w\\ y\\ z\end{bmatrix},\begin{bmatrix}w\\ y\\ z\end{bmatrix}\geq{}0\right\}.$ (9)
The conditions of 1 then provide conditions for solving the optimization in ($i$) whenever:
  1. 1)
    For all $\left(x,u\right)\in\mathcal{P}$, there exists a $v$ such that $\left(Ax+Bu,v\right)\in\mathcal{P}$.
  2. 2)
    $\left(s,r\right)\in\mathcal{P}^{*}$.
  3. 3)
    There exists a suitable function $\phi$.
It is readily checked that 1) amounts to requiring that $A\geq{}\left|{EB}\right|$. For 2) and 3) it is convenient to appeal to 2. Condition ($ii$) of this statement postulates that the set
$\mathcal{C}_{\mu}=\left\{\lambda:\left(\lambda,\mu\right)\in\mathcal{P}^{*}\right\}$
has a minimum element. From eq. 9 we see that for any $\mu$, $y$ and $z$ can be chosen according to
$y=\mu_{-}+v,\;z=\mu_{+}+v,$
where $v\geq{}0$, and the $\mu_{-}$ and $\mu_{+}$ denote the positive and negative parts of the vector $\mu$ respectively. It then follows that $\lambda\in\mathcal{C}_{\mu}$ if and only if
$\lambda=w+E^{\top}\left(\left|{\mu}\right|+v\right),$
where $w\geq{}0$. The minimum element is therefore given by $\lambda^{*}=E^{\top}\left|{\mu}\right|$. This implies that for (2) we require that $s\geq{}E^{\top}\left|{r}\right|$, and also that
$\phi\left(\mu\right)=-E^{\top}\left|{\mu}\right|.$
Condition ($ii$) in 1 then collapses to the existence of a $\lambda_{*}\geq{}0$ such that
$\lambda_{*}=s+A^{\top}\lambda_{*}-E^{\top}\left|{r+B^{\top}\lambda_{*}}\right|,$
which can be checked with linear programming. The convex program in ($iii$) similarly simplifies, this time to the following linear program:
Maximize $\lambda^{\top}x_{0}$
subject to $\begin{bmatrix}s\\ r\end{bmatrix}=\begin{bmatrix}I-A^{\top}&I&E^{\top}&E^{\top}\\ -B^{\top}&0&-I&-I\end{bmatrix}\begin{bmatrix}\lambda\\ w\\ y\\ z\end{bmatrix}\!,\begin{bmatrix}\lambda\\ w\\ y\\ z\end{bmatrix}\!\!\geq{}\!0.$
3.3 Linear Quadratic Regulation with Additional Structure
In this example we will continue our study of the matrix LQR problem outlined at the beginning of section 3.1. However, we now consider the case that the dynamics preserve a cone $\mathcal{P}$ that contains the semi-definite cone. When considering the application of 1, this will implicitly add further constraints to the allowable dynamics. However the upshot is that the optimal control will inherit these structural constraints, meaning that they can be exploited when implementing the optimal control laws. As we will see, this shows for example that the optimal control laws (under suitable definitions) for systems with a circulant structure inherit the same circulant structure [8, 9], with a host of other extensions, for example to systems defined through a generalized frequency variable [10].
The ideas that we are about to present can be significantly generalized, but should serve to illustrate the core ideas. We start with the definition of the cone.
Definition 1.
Let $Q\in\mathbb{R}^{m\times{}m}$ be invertible, and
$T=\begin{bmatrix}Q\otimes{}I_{p}&0\\ 0&Q\otimes{}I_{q}\end{bmatrix}.$
Define
$\mathcal{P}_{Q}=\left\{\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}{}&U_{2}\end{bmatrix}\in\mathbb{S}^{m\left(p+q\right)}:A\left(T\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}{}&U_{2}\end{bmatrix}T^{\top}{}\right)\succeq{}0\right\},$
where $\mathbb{S}^{m\left(p+q\right)}$ denotes the set of ${m\left(p+q\right)}\times{}{m\left(p+q\right)}$ symmetric matrices,
$A\left(\begin{bmatrix}W&X\\ Y&Z\end{bmatrix}\right)=\begin{bmatrix}\operatorname{diag}_{m}\left(W\right)&\operatorname{diag}_{m}\left(X\right)\\ \operatorname{diag}_{m}\left(Y\right)&\operatorname{diag}_{m}\left(Z\right)\end{bmatrix},$
and $\operatorname{diag}_{m}:\mathbb{R}^{mx\times{}my}\rightarrow{}\mathbb{R}^{mx\times{}my}$ denotes the operator that zeros out a given matrix outside a set of $m$ blocks of size $x\times y$ along the diagonal [^3] That is $\operatorname{diag}_{m}\left(M\right)=\begin{bmatrix}M_{11}&0&\cdots{}&0\\ 0&M_{22}&0\\ \vdots{}&0&\ddots{}&0\\ 0&\cdots{}&0&M_{mm}\end{bmatrix},M_{kk}\in\mathbb{R}^{x\times{}y}.$
We now have the following theorem, concerning the matrix LQR problem from section 3.1. The first part of the statement sets up the LQR problem for systems that preserve both the semi-definite cone, and $\mathcal{P}_{Q}$. The important part of the statement is eq. 10. In words, this equation shows that whenever the LQR problem is well defined, the optimal control can be computed through a state-feedback written in terms of matrices structured by the dual cone $\mathcal{P}_{Q}$, revealing the extra structure in the control law.
Theorem 3.
Let $\mathcal{P}_{Q}$ be as in 1, and suppose that $F$, $G$, $S$, $R_{1}$ and $R_{2}$ are such that:
  1. (a)
    $\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\in\mathcal{P}_{Q}$ implies that for some $V_{1}$ and $V_{2}$,
    $\begin{bmatrix}FXF^{\top}+FU_{1}G^{\top}+GU_{1}^{\top}F^{\top}+GU_{2}G^{\top}&V_{1}\\ V_{1}^{\top}&V_{2}\end{bmatrix}\in\mathcal{P}_{Q}.$
  2. (b)
    $\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}&R_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}.$
Then whenever the optimization
Minimize $\sum_{t=0}^{\infty}\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)$
over $\{U_{1}(t),U_{2}(t)\}_{t=0}^{\infty},$
subject to $X(t+1)=\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix},$
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\succeq{}0,X(0)\succeq{}0$
has a finite value, there exists a
$\begin{bmatrix}\Lambda{}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
such that the input sequence $\{\overline{U}_{1}(t),\overline{U}_{2}(t)\}_{t=0}^{\infty}$ that minimizes the cost satisfies
$\begin{bmatrix}\Lambda{}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\begin{bmatrix}X(t)&\overline{U}_{1}(t)\\ \overline{U}_{1}(t)^{\top}&\overline{U}_{2}(t)\end{bmatrix}=0$ (10)
for all $t$.
Proof.
See Appendix A. ∎
To illustrate the theorem, suppose that the matrices $F$ and $G$ have a block-wise circulant structure. That is,
$F=\begin{bmatrix}f_{1}\bar{F}&f_{2}\bar{F}&\cdots{}&f_{m}\bar{F}\\ f_{m}\bar{F}&f_{1}\bar{F}&\cdots{}&f_{m-1}\bar{F}\\ \vdots{}&\vdots{}&\ddots{}&\vdots{}\\ f_{2}\bar{F}&f_{3}\bar{F}&\cdots{}&f_{1}\bar{F}\end{bmatrix},$
with a similar expression for $G$. We may write this compactly according to
$F=F_{\text{sd}}\otimes{}\bar{F}\;\;\text{and}\;\;G=G_{\text{sd}}\otimes{}\bar{G},$ (11)
where $\bar{F}\in\mathbb{R}^{p\times{}p}$ and $\bar{G}\in\mathbb{R}^{p\times{}q}$, and
$\!F_{\text{sd}}\!=\!\!\begin{bmatrix}f_{1}&f_{2}&\cdots{}&f_{m}\\ f_{m}&f_{1}&\cdots{}&f_{m-1}\\ \vdots{}&\vdots{}&\ddots{}&\vdots{}\\ f_{2}&f_{3}&\cdots{}&f_{1}\end{bmatrix}\!\!,G_{\text{sd}}\!=\!\!\begin{bmatrix}g_{1}&g_{2}&\cdots{}&g_{m}\\ g_{m}&g_{1}&\cdots{}&g_{m-1}\\ \vdots{}&\vdots{}&\ddots{}&\vdots{}\\ g_{2}&g_{3}&\cdots{}&g_{1}\end{bmatrix}\!\!.$
Critically for our purposes, the matrices $F_{\text{sd}}$ and $G_{\text{sd}}$ are simultaneously diagonalizable. That is, there exists an invertible matrix $Q$ such that both
$Q^{-1}F_{\text{sd}}Q\;\;\text{and}\;\;Q^{-1}G_{\text{sd}}Q$
are diagonal. This implies that
$\left(Q^{-1}\otimes{}I_{p}\right)F\left(Q\otimes{}I_{p}\right)=\operatorname{diag}_{m}\left(\left(Q^{-1}\otimes{}I_{p}\right)F\left(Q\otimes{}I_{p}\right)\right)$
and
$\left(Q^{-1}\otimes{}I_{p}\right)G\left(Q\otimes{}I_{q}\right)=\operatorname{diag}_{m}\left(\left(Q^{-1}\otimes{}I_{p}\right)G\left(Q\otimes{}I_{q}\right)\right).$
It follows from these expressions that for such matrices, 3(a) holds. As is shown in 1 in the Appendix A,
$\mathcal{P}_{Q}^{*}=\left\{T^{\top}A\left(Y\right)T:Y\succeq{}0\right\},$
where $A$ and $T$ are as in 1. That is, the dual cone consists of the positive semi-definite matrices with blocks that have a similar circulant structure to the matrices $F$ and $G$. 3 then states that provided the optimization criteria is chosen appropriately, the matrix that encodes the optimal control law eq. 10 will have the same block-wise circulant structure. Interestingly this cone based treatment makes no use of any orthogonality properties of the matrix $Q$ allowing similar insights to be derived for general dynamics that take the form of eq. 11, see for example [10].
4 Conclusions
We have proved a general result on optimal control on positive cones, covering standard linear quadratic control and corresponding results for positive systems as special cases. Of course, many other special cases can be derived by defining other relevant cones, or combinations of the ones above. Hopefully, this will in the future form the basis for a powerful toolbox combining the versatility of classical linear quadratic control with the rich scalability properties of positive systems and shortest path problems in networks. Preliminary steps in this direction have already been taken in [11] and [12].
Appendix A Proof of 3
The result will follow from an application of 1, based on the following simple properties of $\mathcal{P}_{Q}$.
Lemma 1.
Let $\mathcal{P}_{Q}$ be as in 1. Then:
  1. ($i$)
    If $P\succeq{}0$, then $P\in\mathcal{P}_{Q}$.
  2. ($ii$)
    $\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*}$
    only if
    $\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*},$
    where $M_{2}\in\mathbb{R}^{mq\times{}mq}$.
Proof.
Claim ($i$) follows immediately from the implication
$Y\succeq{}0\implies{}A\left(Y\right)\succeq{}0.$
For ($ii$), we start by developing a formula for the dual cone. Let
$B\left(X\right)=A\left(TXT^{\top}\right).$
By Farkas’ lemma,
$\mathcal{P}^{*}=\left\{B^{*}\left(Y\right):Y\succeq{}0\right\}.$
Since for any symmetric matrices $C$ and $D$, $\operatorname{tr}\left(CD\right)=\operatorname{vec}\left(C\right)^{\top}\operatorname{vec}\left(D\right)$, we see that
$\operatorname{tr}\left(B\left(X\right)Y\right)$ $=\operatorname{tr}\left(A\left(TXT^{\top}{}\right)A\left(Y\right)\right)$ (12)
$=\operatorname{tr}\left(TXT^{\top}A\left(Y\right)\right)$
$=\operatorname{tr}\left(XT^{\top}A\left(Y\right)T\right).$
This shows that
$B^{*}\left(Y\right)=T^{\top}A\left(Y\right)T.$
It then follows that if
$\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*},$
then there exists a
$\begin{bmatrix}Y_{1}&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}=\begin{bmatrix}\operatorname{diag}_{m}\left(Y_{1}\right)&\operatorname{diag}_{m}\left(Y_{2}\right)\\ \operatorname{diag}\left(Y_{2}^{\top}\right)&\operatorname{diag}_{m}\left(Y_{3}\right)\end{bmatrix}\succeq{}0$
such that
$\begin{bmatrix}0&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}0&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}T.$
It is then straightforward to show that
$\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}Y_{2}Y_{3}^{+}Y_{2}^{\top}&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}T,$
which is in $\mathcal{P}^{*}$ as required. ∎
We now prove 3.
Proof.
Observe that by 1($i$), the optimization in 3 is lower bounded by the relaxed optimization problem
Minimize $\sum_{t=0}^{\infty}\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)$
over $\{U_{1}(t),U_{2}(t)\}_{t=0}^{\infty},$
subject to $X(t+1)=\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix},$
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\in\mathcal{P}_{Q},X(0)\succeq{}0.$
An analogous argument to that made in section 3.1 and 1($ii$) then shows that the set
$\mathcal{C}_{\mu}=\left\{\Lambda:\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}\right\}$
is non-empty, it has a minimum element. Therefore under (a) and (b) in the theorem statement and by 2, the relaxed optimization problem meets all the conditions of 1, from which it soon follows that the relaxed optimization has a finite value if and only if there exists a
$\begin{bmatrix}\Lambda^{*}&0\\ 0&0\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
such that
$\Lambda^{*}=S+F^{\top}\Lambda^{*}F-\left(F^{\top}\Lambda^{*}G+R_{1}\right)\left(R_{2}+G^{\top}{}\Lambda^{*}G\right)^{+}\left(G^{\top}\Lambda^{*}F+R_{1}^{\top}{}\right).$
Again by 1, the elements of the optimal input sequence $\overline{U}_{1}(t)$ and $\overline{U}_{2}(t)$ are the minimizers of the following optimization problem
Minimize $\operatorname{tr}\left(\begin{bmatrix}0&R_{1}+F^{\top}\Lambda^{*}G\\ R_{1}^{\top}+G^{\top}\Lambda^{*}F&R_{2}+G^{\top}\Lambda^{*}G\end{bmatrix}\begin{bmatrix}0&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\right)$
over $U_{1},U_{2},$
subject to $\begin{bmatrix}X(t)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\in\mathcal{P}_{Q}.$
The result will then follow if
$\begin{bmatrix}X(t)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\succeq{}0,$
since then all inputs to the relaxed problem are feasible also for the unrelaxed problem. It follows from an argument analogous to that in eq. 12, and from the assumption that $X(0)\succeq{}0$, that we may replace the constraint in the above with
$\begin{bmatrix}X(0)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\succeq{}0$ (13)
without affecting the optimal value. Just as for the standard LQR problem it then follows from complementary slackness that there exists a $K$ such that
$\begin{bmatrix}I\\ K^{\top}\end{bmatrix}\Lambda^{*}\begin{bmatrix}I&K\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
and
$\begin{bmatrix}I\\ K^{\top}\end{bmatrix}\Lambda^{*}\begin{bmatrix}I&K\end{bmatrix}\begin{bmatrix}X(0)&\overline{U}_{1}\\ \overline{U}_{1}^{\top}&\overline{U}_{2}\end{bmatrix}=0.$ (14)
Applying this control results in $X(1)\succeq{}0$, and rerunning this argument for all $t$ then shows that eqs. 13 and 14 hold for all $t$, not just $t=0$. Hence the optimal actions for the relaxed problem are feasible also for the unrelaxed problem, from which the result follows. ∎

### v2

@@
 The third special case corresponds to linear quadratic control with additional structure, such as spatial invariance.
+Notation
+A cone $\mathcal{P}\subseteq{}\mathbb{R}^{n}$ is said to be proper if it is convex, closed, pointed and has a non-empty interior.
+The dual cone of a proper cone $\mathcal{P}$ is also proper, and given by
+$\mathcal{P}^{*}=\left\{\lambda{}:\lambda{}^{\top}{}x\geq{}0\text{ for all }x\in\mathcal{P}\right\}.$
+All further non-standard notation will be introduced as it is used.
 1 Introduction
-The theory of optimal control has been a cornerstone of control theory ever since the early work by Bellman [1].
-Under standard assumptions, a discrete time optimal control problem with non-negative cost
-Minimize $\sum_{t=0}^{\infty}g(x(t),u(t))$
-subject to $x(t+1)=f(x(t),u(t)),\quad x(0)=x_{0}$
-$u(t)\in U(x(t))$
-is solvable if and only if the Bellman equation
-$J^{*}(x)$ $=\min_{u\in U(x)}\left[g(x,u)+J^{*}(f(x,u))\right]$
-has a non-negative solution $J^{*}$ which is finite for all $x$. The minimizing $u$ gives an optimal control policy. See for example [2, section 3.1].
-Of particular importance are the rare situations when the Bellman equation has an explicit solution.
-The most well known case is probably linear dynamics and quadratic cost, introduced in the pioneering work of [3].
+This paper is concerned with optimal control of the form
+Minimize $\sum_{t=0}^{\infty}[s^{\top}x(t)+r^{\top}u(t)]\text{ over }\{u(t)\}_{t=0}^{\infty}$ (1)
+subject to $x(t+1)=Ax(t)+Bu(t)$
+$(x(t),u(t))\in\mathcal{P},\quad x(0)=x_{0}.$
+where $\mathcal{P}$ is a proper cone in $\mathbb{R}^{n}$. Assumptions are made to make sure that the objective function is non-negative and thus finite if and only if the Bellman equation
+$J^{*}(x)$ $=\min_{(x(t),u(t))\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+J^{*}(Ax+Bu)\right]$
+has a non-negative solution $J^{*}$. The minimizing $u$ gives an optimal control policy. See for example [1, section 3.1].
+The setting above turns out be one of the rare situations when the Bellman equation has an explicit solution.
+Otherwise, the most well known case from the literature is probably linear dynamics and quadratic cost, introduced in the pioneering work of [2].
 Another case of fundamental importance is when both $x$ and $u$ are restricted to finite sets.
 This setting leads to shortest path problems on graphs, with an extensive literature dating back to the work of [3].
-Recently, a third class of problems with explicit solutions to the Bellman equation was introduced in the context of linear systems on the positive orthant in $\mathbb{R}^{n}$, i.e.
-so called positive systems, see [5, 6, 7].
-The purpose of this note is to show that the previous instances where the Bellman equation can be solved explicitly are in fact special cases of a more general class of optimal control problems stated in terms of positive cones.
+Recently, a third class of problems with explicit solutions to the Bellman equation was introduced in the context of linear systems on the positive orthant in $\mathbb{R}^{n}$ (the so called positive systems), see [4, 5].
+A closely related case with bilinear dynamics is treated in [6].
+The purpose of this note is to show that several previous instances where the Bellman equation can be solved explicitly are in fact special cases of a more general class of optimal control problems stated in terms of positive cones.
 2 Main Result
 Theorem 1.
-Let $\mathcal{P}$ be a positive cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$.
-Given $(s,r)\in\mathcal{P}^{*}$ and $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, suppose that for every $(x,u)\in\mathcal{P}$ there exists $v$ such that $(Ax+Bu,v)\in\mathcal{P}$.
+Let $\mathcal{P}$ be a proper cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$.
+Given $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, suppose that for every $(x,u)\in\mathcal{P}$ there exists $v$ such that $(Ax+Bu,v)\in\mathcal{P}$.
+Let $(s,r)$ be an interior point of $\mathcal{P}^{*}$.
 Moreover, assume the existence of $\phi:\mathbb{R}^{m}\to\mathbb{R}^{n}$ such that
 $\phi(\mu)^{\top}x$ $=\min_{(x,u)\in\mathcal{P}}\mu^{\top}u$ (2)
-for all $(x,v)\in\mathcal{P}$, $(\lambda,\mu)\in\mathcal{P}^{*}$. Then the following three statements are equivalent:
+for all $(x,v)\in\mathcal{P}$, $(\eta,\mu)\in\mathcal{P}^{*}$. Then the following three statements are equivalent:
   • ($i$)
-    The following optimal control problem has a finite value for every $(x_{0},u_{0})\in\mathcal{P}$:
-    Minimize $\sum_{t=0}^{\infty}[s^{\top}x(t)+r^{\top}u(t)]\text{ over }\{u(t)\}_{t=0}^{\infty}$
-    subject to $x(t+1)=Ax(t)+Bu(t)$
-    $(x(t),u(t))\in\mathcal{P},\quad x(0)=x_{0}.$
+    The problem eq. 1 has a finite value for all $(x_{0},u_{0})\in\mathcal{P}$.
   • ($ii$)
-    There exists $(\lambda_{*},0)\in\mathcal{P}^{*}$ such that
-    $\lambda_{*}=s+A^{\top}\lambda_{*}+\phi(r+B^{\top}\lambda_{*}).$ (2)
+    There exists a $(\lambda_{*},0)\in\mathcal{P}^{*}$ such that
+    $\lambda_{*}=s+A^{\top}\lambda_{*}+\phi(r+B^{\top}\lambda_{*}).$
   • ($iii$)
@@
 Proof.
-The minimization in ($i$) is a special case of the general formulation in the introduction, where $f(x,u)=Ax+Bu$ and $g(x,u)=s^{\top}x+r^{\top}u$.
-Define the sequence $\{\lambda_{k}\}_{k=0}^{\infty}$ recursively by
-$\lambda_{k+1}$ $:=s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k}),$ $\lambda_{0}$ $:=0.$
-If $J_{k}(x)=\lambda_{k}^{\top}x$, then
-$\min_{u\in U(x)}\left[g(x,u)+J_{k}(f(x,u))\right]$
+Given $(\lambda_{0},0)\in\mathcal{P}^{*}$, consider the Bellman recursion
+$J_{k+1}(x)$ $=\min_{x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+J_{k}(Ax+Bu)\right],$
+$J_{0}(x)$ $=\lambda_{0}^{\top}x$
+and the sequence $\{\lambda_{k}\}_{k=0}^{\infty}$ generated by
+$\lambda_{k+1}$ $=s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k}).$ (3)
+We will prove by induction that this implies $J_{k}(x)=\lambda_{k}^{\top}x$ for all $k$: Assume for some $k\geq 0$ that $J_{k}(x)=\lambda_{k}^{\top}x$ and $(\lambda_{k},0)\in\mathcal{P}^{*}$.
+Let $(\eta_{k},\mu_{k}):=(s+A^{\top}\lambda_{k},r+B^{\top}\lambda_{k})$.
+Then
+$\eta_{k}^{\top}x+\mu_{k}^{\top}u$ $=s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)\geq 0$
+for all $(x,u)\in\mathcal{P}$, so $(\eta_{k},\mu_{k})\in\mathcal{P}^{*}$. Hence
+$J_{k+1}(x)$ $=\min_{x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+J_{k}(Ax+Bu)\right]$
 $=\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)\right]$
-$=s^{\top}x+\lambda_{k}^{\top}Ax+\min_{(x,u)\in\mathcal{P}}\left[(r+B^{\top}\lambda_{k})^{\top}u\right]$
-$=\left[s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k})\right]^{\top}x$
-$=\lambda_{k+1}^{\top}x.$
-Hence, value iteration
-$J_{k+1}(x)$ $=\min_{u\in U(x)}\left[g(x,u)+J_{k}(f(x,u))\right],$ $J_{0}$ $=0$
-gives a sequence $0=J_{0}\leq J_{1}\leq J_{2}\leq\ldots$ with $J_{k}(x)=\lambda_{k}^{\top}x$ for all $k$.
-In particular $(\lambda_{k},0)\in\mathcal{P}_{*}$ for all $k$.
-If ($i$) holds, then the sequence has an upper limit $J^{*}(x)=\lambda_{*}^{\top}x$ with $\lambda_{*}$ satisfying the condition in ($ii$).
-Conversely, if ($ii$) holds, then $(\lambda_{*}-\lambda_{k},0)\in\mathcal{P}_{*}$ for all $k$, so the sequence $J_{0}\leq J_{1}\leq J_{2}\leq\ldots$ has an upper limit and ($i$) holds too.
-This proves the equivalence of ($i$) and ($ii$), as well as the formula $\lambda_{*}^{\top}x_{0}$ for the minimal value in ($i$).
-The optimal controller is $u(t)=\lambda(x(t)$, where
-$\lambda(x)$ $=\arg\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{*}^{\top}(Ax+Bu)\right]$
-$=\arg\min_{(x,u)\in\mathcal{P}}\left[(r+B^{\top}\lambda_{*})^{\top}u\right].$
-It remains to prove equivalence betwen ($ii$) and ($iii$).
-First, by definition of $\phi$, the optimization constraint in ($iii$) implies that
-$\lambda^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda^{\top}(Ax+Bu)$ (3)
-for all $(x,u)\in\mathcal{P}$.
-It is a well known fact that solutions to the Bellman inequality are lower bounds on the optimal cost.
-Hence, if ($ii$) holds, the maximum for the convex program must be attained at the solution of eq.
-2, so also ($iii$) holds.
-Conversely, suppose that ($iii$) holds with the maximizing argument $\lambda_{k}$.
-Define
-$\lambda_{k+1}$ $:=s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k}).$
-It follows from eq. 3 that
-$\lambda_{k}^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)$
-$\lambda_{k}^{\top}x$ $\leq s^{\top}x+\lambda_{k}^{\top}Ax+\phi(r+B^{\top}\lambda_{k})^{\top}x$
-for all $(x,u)\in\mathcal{P}$ and therefore
-$(\lambda_{k+1}-\lambda_{k},0)$ $=(s+A^{\top}\lambda_{k}+\phi(r+B^{\top}\lambda_{k})-\lambda_{k},0)\in\mathcal{P}_{*}.$
-It follows that
-$\lambda_{k+1}^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda_{k+1}^{\top}(Ax+Bu)$
-Hence also $\lambda_{k+1}$ satisfies the constraints of the optimization in ($iii$), so $\lambda_{k+1}=\lambda_{k}$.
-Otherwise $\lambda_{k}$ would not be maximizing.
-This means that $\lambda_{k}$ solves eq.
-2 and also ($ii$) holds.
+$=s^{\top}x+\lambda_{k}^{\top}Ax+\min_{(x,u)\in\mathcal{P}}\left[\mu_{k}^{\top}u\right]$
+$=\left[s+A^{\top}\lambda_{k}+\phi(\mu_{k})\right]^{\top}x=\lambda_{k+1}^{\top}x$
+and $\lambda_{k+1}^{\top}x\geq 0$, so $(\lambda_{k+1},0)\in\mathcal{P}^{*}$ and the induction assumption holds also for $k+1$.
+The assumption is trivial for $k=0$, so it follows by induction for $k=1,2,3,\ldots$
+($ii$) $\Longleftrightarrow$ ($i$): If ($i$) holds, then the Bellman recursion from $J_{0}=0$ gives $0=J_{0}\leq J_{1}\leq J_{2}\leq\ldots$ with upper limit $\lambda_{*}^{\top}x$ where $\lambda_{*}$ satisfies the condition in ($ii$).
+Conversely, if ($ii$), then $\lambda_{0}=0$ gives $(\lambda_{*}-\lambda_{k},0)\in\mathcal{P}^{*}$ for all $k$.
+This is trivial for $k=0$ and follows by induction for all other $k$ since
+$\lambda_{k+1}^{\top}x$ $=\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu)\right]$
+$\leq\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{*}^{\top}(Ax+Bu)\right]=\lambda_{*}^{\top}x.$
+Hence the $J_{k}$-sequence has an upper limit and ($i$) holds too.
+This proves the implication from ($ii$) to ($i$), as well as the formula $\lambda_{*}^{\top}x_{0}$ for the minimal value in ($i$).
+Given $x$, the optimal control is $u$ is obtained by the minimization
+$\arg\min_{(x,u)\in\mathcal{P}}\left[s^{\top}x+r^{\top}u+\lambda_{*}^{\top}(Ax+Bu)\right]$
+or equivalently $\arg\min_{(x,u)\in\mathcal{P}}\left[(r+B^{\top}\lambda_{*})^{\top}u\right]$.
+($ii$$\implies$($iii$): Consider $\lambda_{*}$ satisfying ($ii$) and let $\lambda$ be a feasible point for the optimization problem in ($iii$). In particular,
+$\lambda^{\top}x$ $\leq s^{\top}x+r^{\top}u+\lambda^{\top}(Ax+Bu)$ for ($\text{for }(x,u)\in\mathcal{P}.$ (4)
+The assumption that $(s,r)$ is an interior point of $\mathcal{P}^{*}$ gives existence of $\gamma,\delta>0$ such that
+$\lambda_{*}^{\top}(Ax+Bu)$ $\leq\gamma(s^{\top}x+r^{\top}u)$
+$\lambda^{\top}x$ $\leq\delta s^{\top}x\leq\delta\lambda_{*}^{\top}x$ for ($\text{for }(x,u)\in\mathcal{P}.$
+Let $\{\lambda_{k}\}_{k=0}^{\infty}$ be the sequence generated by eq.
+3 starting from $\lambda_{0}=\lambda$.
+Combination of eq.
+3 and eq.
+4 gives $(\lambda_{k}-\lambda,0)\in\mathcal{P}^{*}$ for all $k$.
+Moreover, [7, Proposition 2] implies that
+00 0 $\leq\lambda_{k}^{\top}x\leq\left[1+\frac{\delta-1}{(1+\gamma^{-1})^{k}}\right]\lambda_{*}^{\top}x$ for all $\text{for all }k,$
+so in the limit $(\lambda_{*}-\lambda,0)\in\mathcal{P}$ and $\lambda_{*}^{\top}x$ an upper bound for the value of the convex program in ($iii$).
+($iii$$\implies$($ii$): Consider the sequence defined by eq. 3 initialized with $\lambda_{0}=0$. Then
+$\lambda_{k}^{\top}x$ $\leq\lambda_{k+1}^{\top}x=s^{\top}x+r^{\top}u+\lambda_{k}^{\top}(Ax+Bu),$
+so $\lambda_{k}$ is a feasible point for the optimization problem in ($iii$).
+Hence ($iii$) implies that the sequence $\lambda_{1}^{\top}x\leq\lambda_{2}^{\top}x\leq\ldots$ is bounded for every $x$.
+Therefore the limit $\lambda_{*}=\lim_{k\to\infty}\lambda_{k}$ exists and satisfies ($ii$).
 The proof is complete.
@@
 It is natural then to ask when such a $\phi$ exists.
-The following result shows that when the cone $\mathcal{P}$ is proper, this can be established in a rather direct way through a property of the dual cone $\mathcal{P}^{*}$.
+The following result shows that this can be established in a rather direct way through a property of the dual cone $\mathcal{P}^{*}$.
 We will use this result in the next section when we connect 1 to some well studied optimal control problems.
 Theorem 2.
-Let $\mathcal{P}$ be a proper cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$, $\lambda^{*}\in\mathbb{R}^{n}$, and $\mu\in\mathbb{R}^{m}$.
-Suppose that there exists a $\lambda\in\mathbb{R}^{n}$ such that
-$\left(\lambda,\mu\right)\in\operatorname{rel}\operatorname{int}\mathcal{P}^{*}.$ (4)
-Then the following statements are equivalent:
+Let $\mathcal{P}$ be a proper cone in $\mathbb{R}^{n+m}$ with dual cone $\mathcal{P}^{*}$. For any $\mu\in\mathbb{R}^{m}$ such that the set
+$\mathcal{C}_{\mu}=\left\{\lambda:\left(\lambda,\mu\right)\in\mathcal{P}^{*}\right\}$ (5)
+has a non-empty interior, the following are equivalent:
   1. ($i$)
-    $-\lambda^{*\top{}}x=\inf_{\left(x,u\right)\in\mathcal{P}}\mu^{\top}u.$ (5)
+    The function
+    $x\mapsto\inf_{\left(x,u\right)\in\mathcal{P}}\mu^{\top}u$ (6)
+    with domain $\left\{x:\left(x,v\right)\in\mathcal{P}\text{ for some }v\right\}$ is a linear form (that is, the function above takes the form $f(x)=\phi^{\top}x$ for some $\phi\in\mathbb{R}^{n}$).
   2. ($ii$)
-    $\left(\lambda^{*},\mu\right)\in\mathcal{P}^{*},\;\text{and for all}\;\left(\lambda,\mu\right)\in\mathcal{P}^{*},\;\left(\lambda-\lambda^{*},0\right)\in\mathcal{P}^{*}.$
+    The set $\mathcal{C}_{\mu}$ has a minimum element (that is, there exists a $\bar{\lambda}\in\mathcal{C}_{\mu}$ such that $\lambda\in\mathcal{C}_{\mu}\implies\left(\lambda-\bar{\lambda},0\right)\in\mathcal{P}^{*}$).
+Furthermore, if $\bar{\lambda}$ is the minimum element in ($ii$), then the linear form in ($i$) is $f(x)=-\bar{\lambda}^{\top}x$.
 Proof.
-($i$$\implies$($ii$): With a view to generating a contradiction, suppose that under the hypothesis of ($i$),
-$\left(\lambda^{*},\mu\right)\notin\mathcal{P}^{*}.$
+($i$$\implies$($ii$): Under the hypothesis of ($i$), there exists a $\bar{\lambda}$ such that the function in eq.
+6 takes the form $f\left(x\right)=-\bar{\lambda}^{\top}{}x$.
+We will first establish that $\left(\bar{\lambda},\mu\right)\in\mathcal{P}^{*}$.
+Assume the converse.
 There therefore exists an
-$\left(\bar{x},\bar{u}\right)\in\mathcal{P}\;\text{such that}\;\begin{bmatrix}\lambda^{*}\\ \mu\end{bmatrix}^{\top}\begin{bmatrix}\bar{x}\\ \bar{u}\end{bmatrix}<0,$
-which implies that $-\lambda^{*\top{}}\bar{x}>\mu^{\top}\bar{u}$.
+$\left(\bar{x},\bar{u}\right)\in\mathcal{P}\;\text{such that}\;\begin{bmatrix}\bar{\lambda}\\ \mu\end{bmatrix}^{\top}\begin{bmatrix}\bar{x}\\ \bar{u}\end{bmatrix}<0,$
+which implies that $-\bar{\lambda}^{\top}\bar{x}>\mu^{\top}\bar{u}$.
 However from eq.
-5 we see that $-\lambda^{*\top{}}\bar{x}\leq{}\mu^{\top}\bar{u}$, which is a contradiction.
-This proves the first claim in ($ii$).
-Now, again with a view to generating a contradiction, suppose that
-$\left(\lambda,\mu\right)\in\mathcal{P}^{*}\;\text{and}\;\left(\lambda-\lambda^{*},0\right)\notin\mathcal{P}^{*}.$
-Therefore there exists an
-$\left(\tilde{x},\tilde{u}\right)\in\mathcal{P}\;\text{such that}\;\begin{bmatrix}\lambda-\lambda^{*}\\ 0\end{bmatrix}^{\top}\begin{bmatrix}\tilde{x}\\ \tilde{u}\end{bmatrix}<0,$
-which implies that $\lambda^{*\top{}}\tilde{x}>\lambda^{\top}\tilde{x}$. The dual of eq. 5 reads
+6 we see that $-\bar{\lambda}^{\top{}}\bar{x}\leq{}\mu^{\top}\bar{u}$, which is a contradiction.
+We will now show that $\bar{\lambda}$ is the minimum element of $\mathcal{C}_{\mu}$.
+The optimization in eq.
+6 is a cone program on standard form, with dual program (see for example [8, p.266])
 $\sup_{\left(\lambda,\mu\right)\in\mathcal{P}^{*}}-\lambda^{\top}x.$ (7)
-The solution to eq. 6 is lower bounded by $-\lambda^{\top}\tilde{x}$. However by weak duality, the solution to eq. 6 is upper bounded by $-\lambda^{*\top{}}\tilde{x}$. Therefore
-$\left(\lambda,\mu\right)\in\mathcal{P}^{*}\implies\lambda^{*\top{}}\tilde{x}\leq{}\lambda^{\top}\tilde{x}.$
-Again this a contradiction, proving the claim.
-($ii$$\implies$($i$): Since $\mathcal{P}$ is a proper cone, $\mathcal{P}^{**}=\mathcal{P}$. Therefore for all $\left(x,u\right)\in\mathcal{P}$,
-$\begin{bmatrix}\lambda-\lambda^{*}\\ 0\end{bmatrix}^{\top}\begin{bmatrix}x\\ u\end{bmatrix}\geq{}0,\implies-\left(\lambda-\lambda^{*}\right)^{\top}x\leq{}0.$
+By weak duality this implies that
+$\inf_{\left(x,u\right)\in\mathcal{P}}\mu^{\top}u\geq{}\sup_{\left(\lambda,\mu\right)\in\mathcal{P}^{*}}-\lambda^{\top}x,$
+and so for any $\left(x,u\right)\in\mathcal{P}$,
+$\left(\lambda,\mu\right)\in\mathcal{P}^{*}\implies\bar{\lambda}^{\top{}}x\leq{}\lambda^{\top}x.$ (8)
+This implies that $\left(\lambda-\bar{\lambda},0\right)\in\mathcal{P}^{*}$, which proves the claim.
+($ii$$\implies$($i$): For all $\left(x,u\right)\in\mathcal{P}$,
+$\begin{bmatrix}\lambda-\bar{\lambda}\\ 0\end{bmatrix}^{\top}\begin{bmatrix}x\\ u\end{bmatrix}\geq{}0,\implies-\left(\lambda-\bar{\lambda}\right)^{\top}x\leq{}0.$
 Therefore for any such $\lambda$,
-$-\lambda^{\top}x=-\left(\lambda-\lambda^{*}\right)^{\top}x-\lambda^{*\top{}}x\leq{}-\lambda^{*\top{}}x.$
-It then follows that the solution to eq.
-6 equals $-\lambda^{*\top{}}x$.
-Condition eq.
-4 states that eq.
-6 is strictly feasible.
-This implies that strong duality holds, which proves eq.
-5 as required.
-∎
+$-\lambda^{\top}x=-\left(\lambda-\bar{\lambda}\right)^{\top}x-\bar{\lambda}^{\top{}}x\leq{}-\bar{\lambda}^{\top{}}x.$
+It then follows that the solution to eq. 7 equals $-\bar{\lambda}^{\top{}}x$. Condition eq. 5 ensures that eq. 7 is strictly feasible. Therefore strong duality holds, which proves eq. 6 as required. ∎
 3 Examples
@@
 In the above, the variables are the matrices $X(t)$, $U_{1}(t)$ and $U_{2}(t)$. The problem data is the matrices $S$, $R_{1}$, $R_{2}$, $F$ and $G$, where we additionally require that
-$\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\succeq{}0.$ (7)
+$\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\succ{}0.$ (9)
 Although it may not be immediately clear at first sight, this is an optimization of precisely the type considered in 1($i$) where $\mathcal{P}$ is the positive semi-definite cone (which is positive, proper and self-dual).
@@
 $\mathcal{C}_{\mu}=\left\{\Lambda:\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\succeq{}0\right\},$
-having a minimum element (that is, an element $\Lambda^{*}\in\mathcal{C}_{\mu}$ such that $\Lambda^{*}\preceq{}\Lambda$ for all $\Lambda\in\mathcal{C}_{\mu}$).
-For the semi-definite cone whenever $\mathcal{C}_{\mu}$ is non-empty, this element exists, and is given by $\Lambda^{*}=M_{1}M_{2}^{+}M_{1}^{\top}{}$.
-Therefore
+having a minimum element. For the semi-definite cone whenever $\mathcal{C}_{\mu}$ is non-empty, this element exists, and is given by $\overline{\Lambda}=M_{1}M_{2}^{+}M_{1}^{\top}{}$. Therefore
 $\phi\left(\mu\right)\longleftrightarrow{}-M_{1}M_{2}^{+}M_{1}^{\top}{}.$
@@
 Substituting in for all found correspondances, we see that in our matrix notation [^2] It is perhaps worth noting that from basic properties of adjoints, it follows that $A^{\top}\lambda(t)$ $\longleftrightarrow{}F^{\top}\Lambda(t)F$ $B^{\top}\lambda(t)$ $\longleftrightarrow{}\begin{bmatrix}0&F^{\top}\Lambda(t)G\\ G^{\top}\Lambda(t)F&G^{\top}\Lambda(t)G\end{bmatrix}.$ , the equation $\lambda_{*}=s+A^{\top}\lambda_{*}+\phi(r+B^{\top}\lambda_{*})$ becomes
-$\Lambda^{*}=S+F^{\top}\Lambda^{*}F-\left(F^{\top}\Lambda^{*}G+R_{1}\right)\left(R_{2}+G^{\top}{}\Lambda^{*}G\right)^{+}\left(G^{\top}\Lambda^{*}F+R_{1}^{\top}{}\right).$
-Note that a degree of care has to be taken when reading this formula, since the requirements on the domain of $\phi$ have been left implicit.
-For example, these are guaranteed to hold under the common assumption that $R_{2}\succ{}0$, after which the above collapses to the usual algebraic Riccati equation.
+$\Lambda_{*}$ $=S+F^{\top}\Lambda_{*}F-\left(F^{\top}\Lambda_{*}G+R_{1}\right)\left(R_{2}+G^{\top}{}\Lambda_{*}G\right)^{-1}\left(G^{\top}\Lambda_{*}F+R_{1}^{\top}{}\right).$
+That is, the equation in 1($ii$) becomes the algebraic Riccati equation (the pseudo-inverses become inverses in light of eq.
+9).
 This can then be solved in the usual way whenever the pair $\left(F,G\right)$ is stabilizable.
-1 ($iii$) can be similarly analyzed.
+1($iii$) can be similarly analyzed.
 In this case, the convex program reduces to the semi-definite program
@@
 and also that
-$\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)=$
-$\qquad\qquad\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}^{\top}\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}.$
+$\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)=\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}^{\top}\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}.$
 That is, under eq.
-8, the costs and constraints in the standard LQR problem, and the matrix optimization problem, are identical.
-It then follows (with a little extra work) that the standard LQR problem corresponds to the special case of the matrix optimization when the initial condition $X_{0}$ is rank 1.
+10, the costs and constraints in the standard LQR problem, and the matrix optimization problem, are identical.
+It then follows (with a little extra work) that when the initial condition $X_{0}$ is rank 1, that the studied matrix optimization collapses to the standard LQR problem.
 3.2 A Linear Regulator on a Polyhedral Cone
-In this subsection we consider the application of 1 when $\mathcal{P}$ is the polyhedral cone (which are always positive and proper)
+In this subsection we will recover the results of [4, Theorem 1]. Consider the application of 1 when $\mathcal{P}$ is the polyhedral cone
 $\mathcal{P}=\left\{\begin{bmatrix}x\\ u\end{bmatrix}:\begin{bmatrix}I&0\\ E&-I\\ E&I\end{bmatrix}\begin{bmatrix}x\\ u\end{bmatrix}\geq{}0\right\},$
@@
 $\left|{u}\right|\leq{}Ex.$
-When minimising the cost in 1 ($i$), we are in effect optimizing performance under the constraint that the control keeps the system state positive.
-In what follows will recover the results of [5, Theorem 1].
+When minimizing the cost in 1($i$), we are in effect optimizing performance under the constraint that the control keeps the system state positive.
 First note that the dual cone $\mathcal{P}^{*}$ is also polyhedral, and is given by
@@
     There exists a suitable function $\phi$.
-It is readily checked that 1) amounts to requiring that $A\geq{}\left|{EB}\right|$. For 2) and 3) it is convenient to appeal to 2. Condition ($ii$) of this statement postulates that the set
-$\mathcal{C}_{\mu}=\left\{\lambda:\left(\lambda,\mu\right)\in\mathcal{P}^{*}\right\}$
-has a minimum element. From eq. 9 we see that for any $\mu$, $y$ and $z$ can be chosen according to
+It is readily checked that 1) amounts to requiring that $A\geq{}\left|{EB}\right|$.
+For 2) and 3) it is convenient to appeal to 2.
+From eq.
+11 we see that for any $\mu$, $y$ and $z$ can be chosen according to
 $y=\mu_{-}+v,\;z=\mu_{+}+v,$
@@
 $\lambda=w+E^{\top}\left(\left|{\mu}\right|+v\right),$
-where $w\geq{}0$. The minimum element is therefore given by $\lambda^{*}=E^{\top}\left|{\mu}\right|$. This implies that for (2) we require that $s\geq{}E^{\top}\left|{r}\right|$, and also that
+where $w\geq{}0$.
+The set in 2($ii$) therefore has a minimum element given by $\overline{\lambda}=E^{\top}\left|{\mu}\right|$.
+This implies that for (2) we require that $s>E^{\top}\left|{r}\right|$, and also that
 $\phi\left(\mu\right)=-E^{\top}\left|{\mu}\right|.$
@@
 subject to $\begin{bmatrix}s\\ r\end{bmatrix}=\begin{bmatrix}I-A^{\top}&I&E^{\top}&E^{\top}\\ -B^{\top}&0&-I&-I\end{bmatrix}\begin{bmatrix}\lambda\\ w\\ y\\ z\end{bmatrix}\!,\begin{bmatrix}\lambda\\ w\\ y\\ z\end{bmatrix}\!\!\geq{}\!0.$
-3.3 Linear Quadratic Regulation with Additional Structure
+3.3 A Structured Linear Quadratic Regulation Problem
 In this example we will continue our study of the matrix LQR problem outlined at the beginning of section 3.1.
@@
 and $\operatorname{diag}_{m}:\mathbb{R}^{mx\times{}my}\rightarrow{}\mathbb{R}^{mx\times{}my}$ denotes the operator that zeros out a given matrix outside a set of $m$ blocks of size $x\times y$ along the diagonal [^3] That is $\operatorname{diag}_{m}\left(M\right)=\begin{bmatrix}M_{11}&0&\cdots{}&0\\ 0&M_{22}&0\\ \vdots{}&0&\ddots{}&0\\ 0&\cdots{}&0&M_{mm}\end{bmatrix},M_{kk}\in\mathbb{R}^{x\times{}y}.$
-We now have the following theorem, concerning the matrix LQR problem from section 3.1.
-The first part of the statement sets up the LQR problem for systems that preserve both the semi-definite cone, and $\mathcal{P}_{Q}$.
-The important part of the statement is eq.
-10.
-In words, this equation shows that whenever the LQR problem is well defined, the optimal control can be computed through a state-feedback written in terms of matrices structured by the dual cone $\mathcal{P}_{Q}$, revealing the extra structure in the control law.
-Theorem 3.
-Let $\mathcal{P}_{Q}$ be as in 1, and suppose that $F$, $G$, $S$, $R_{1}$ and $R_{2}$ are such that:
-  1. (a)
-    $\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\in\mathcal{P}_{Q}$ implies that for some $V_{1}$ and $V_{2}$,
-    $\begin{bmatrix}FXF^{\top}+FU_{1}G^{\top}+GU_{1}^{\top}F^{\top}+GU_{2}G^{\top}&V_{1}\\ V_{1}^{\top}&V_{2}\end{bmatrix}\in\mathcal{P}_{Q}.$
-  2. (b)
-    $\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}&R_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}.$
-Then whenever the optimization
-Minimize $\sum_{t=0}^{\infty}\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)$
-over $\{U_{1}(t),U_{2}(t)\}_{t=0}^{\infty},$
-subject to $X(t+1)=\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix},$
-$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\succeq{}0,X(0)\succeq{}0$
-has a finite value, there exists a
-$\begin{bmatrix}\Lambda{}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
-such that the input sequence $\{\overline{U}_{1}(t),\overline{U}_{2}(t)\}_{t=0}^{\infty}$ that minimizes the cost satisfies
-$\begin{bmatrix}\Lambda{}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\begin{bmatrix}X(t)&\overline{U}_{1}(t)\\ \overline{U}_{1}(t)^{\top}&\overline{U}_{2}(t)\end{bmatrix}=0$ (10)
-for all $t$.
+The following lemma establishes the key properties of the cone $\mathcal{P}_{Q}$ that we will require to apply 1.
+The first part of the result gives a formula for the dual cone, and establishes that a suitable function $\phi\left(\mu\right)$ exists (the function is in fact the same as that from section 3.1).
+The final part gives a formula that can be used to compute the optimal control law, and it is from this formula that we can deduce that the optimal control inherits additional structural features as claimed above.
+Lemma 1.
+Let $A$, $T$, $\mathcal{P}_{Q}$ and all the block partitioning of matrices be as in 1. Then
+$\mathcal{P}_{Q}^{*}=\left\{T^{\top}A\left(Y\right)T:Y\succeq{}0\right\}.$
+Furthermore if there exists a $\Lambda$ such that
+$\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*}_{Q},$ (12)
+then for fixed $X$
+$\inf\!\left\{\!\operatorname{tr}\!\left(\!\begin{bmatrix}0&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\!\!\begin{bmatrix}0&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\!\right)\!:\!\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\!\in\mathcal{P}_{Q}\!\right\}$ (13)
+equals $\operatorname{tr}\left(M_{1}M_{2}^{+}M_{1}^{\top}X\right)$. A pair $\left(U_{1},U_{2}\right)$ that achieves the infimum above can be computed through the equation
+$\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}=\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}X\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}^{\top}{}.$ (14)
 Proof.
 See Appendix A. ∎
-To illustrate the theorem, suppose that the matrices $F$ and $G$ have a block-wise circulant structure. That is,
+To understand how to leverage 1, consider again the optimal control problem from section 3.1, but suppose that the matrices $F$ and $G$ have a block-wise circulant structure. That is,
 $F=\begin{bmatrix}f_{1}\bar{F}&f_{2}\bar{F}&\cdots{}&f_{m}\bar{F}\\ f_{m}\bar{F}&f_{1}\bar{F}&\cdots{}&f_{m-1}\bar{F}\\ \vdots{}&\vdots{}&\ddots{}&\vdots{}\\ f_{2}\bar{F}&f_{3}\bar{F}&\cdots{}&f_{1}\bar{F}\end{bmatrix},$
@@
 $\left(Q^{-1}\!\otimes{}I_{p}\right)G\left(Q\otimes{}I_{q}\right)\!=\!\operatorname{diag}_{m}\left(\left(Q^{-1}\!\otimes{}I_{p}\right)G\left(Q\otimes{}I_{q}\right)\right).$
-It follows from these expressions that for such matrices, 3(a) holds. As is shown in 1 in the Appendix A,
-$\mathcal{P}_{Q}^{*}=\left\{T^{\top}A\left(Y\right)T:Y\succeq{}0\right\},$
-where $A$ and $T$ are as in 1.
-That is, the dual cone consists of the positive semi-definite matrices with blocks that have a similar circulant structure to the matrices $F$ and $G$.
-3 then states that provided the optimization criteria is chosen appropriately, the matrix that encodes the optimal control law eq.
-10 will have the same block-wise circulant structure.
+It follows from these expressions that for such matrices,
+$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\in\mathcal{P}_{Q}\implies\begin{bmatrix}X(t+1)&0\\ 0&0\end{bmatrix}\in\mathcal{P}_{Q},$
+where
+$X(t+1)=\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix}.$
+Taking an analogous approach to that in section 3.1, we see that under suitable assumptions on the cost function, 1 can be applied to solve a relaxed version of the standard LQR problem, where the state and inputs can be drawn from $\mathcal{P}_{Q}$.
+Since $\mathcal{P}_{Q}$ contains the positive semi-definite cone, the solution to this problem will always give a lower bound on the achievable cost to the standard LQR problem.
+However more is true.
+In particular we see from the final part of 1 and eq.
+14 in 1 that the optimal inputs to the relaxed problem can be computed through the control law $X\mapsto{}\left(U_{1},U_{2}\right)$ defined implicitly by
+$\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}=\begin{bmatrix}I\\ -K\end{bmatrix}X\begin{bmatrix}I\\ -K\end{bmatrix}^{\top}{},$
+where
+$K=\left(R_{2}+G^{\top}{}\Lambda_{*}G\right)^{-1}\left(G^{\top}\Lambda_{*}F+R_{1}^{\top}{}\right).$
+It follows from these expressions that whenever the initial condition satisfies $X\left(0\right)\succeq{}0$,
+$\begin{bmatrix}X\left(t\right)&U_{1}\left(t\right)\\ U_{1}\left(t\right)^{\top}&U_{2}\left(t\right)\end{bmatrix}\succeq{}0$
+for all $t$. Hence the found controls are also feasible, and hence also optimal, for the standard LQR problem. It is then easily seen from the expression for the dual cone that
+$\left(Q^{-1}\!\otimes{}I_{q}\right)K\left(Q\otimes{}I_{p}\right)\!=\!\operatorname{diag}_{m}\left(\left(Q^{-1}\!\otimes{}I_{q}\right)K\left(Q\otimes{}I_{p}\right)\right).$
+Therefore the optimal control law additionally inherits the block-wise circulant structure present in the problem data.
 Interestingly this cone based treatment makes no use of any orthogonality properties of the matrix $Q$ allowing similar insights to be derived for general dynamics that take the form of eq.
-11, see for example [10].
+15, see for example [11].
 4 Conclusions
@@
 Of course, many other special cases can be derived by defining other relevant cones, or combinations of the ones above.
-Hopefully, this will in the future form the basis for a powerful toolbox combining the versatility of classical linear quadratic control with the rich scalability properties of positive systems and shortest path problems in networks.
-Preliminary steps in this direction have already been taken in [11] and [12].
-Appendix A Proof of 3
-The result will follow from an application of 1, based on the following simple properties of $\mathcal{P}_{Q}$.
-Lemma 1.
-Let $\mathcal{P}_{Q}$ be as in 1. Then:
-  1. ($i$)
-    If $P\succeq{}0$, then $P\in\mathcal{P}_{Q}$.
-  2. ($ii$)
-    $\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*}$
-    only if
-    $\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*},$
-    where $M_{2}\in\mathbb{R}^{mq\times{}mq}$.
+Hopefully, this will in the future form the basis for a powerful toolbox combining the versatility of linear quadratic control with the rich scalability properties of positive systems and shortest path problems in networks.
+Preliminary steps in this direction have been taken in [12] and [13].
+5 Acknowledgment
+The authors are grateful to Yuchao Li for pointing out an error in the original statement of the main result.
+Appendix A Proof of 1
 Proof.
-Claim ($i$) follows immediately from the implication
-$Y\succeq{}0\implies{}A\left(Y\right)\succeq{}0.$
-For ($ii$), we start by developing a formula for the dual cone. Let
-$B\left(X\right)=A\left(TXT^{\top}\right).$
-By Farkas’ lemma,
-$\mathcal{P}^{*}=\left\{B^{*}\left(Y\right):Y\succeq{}0\right\}.$
+We start by developing the formula for the dual cone. Let $B\left(X\right)=A\left(TXT^{\top}\right)$. By Farkas’ lemma,
+$\mathcal{P}^{*}_{Q}=\left\{B^{*}\left(Y\right):Y\succeq{}0\right\}.$
 Since for any symmetric matrices $C$ and $D$, $\operatorname{tr}\left(CD\right)=\operatorname{vec}\left(C\right)^{\top}\operatorname{vec}\left(D\right)$, we see that
-$\operatorname{tr}\left(B\left(X\right)Y\right)$ $=\operatorname{tr}\left(A\left(TXT^{\top}{}\right)A\left(Y\right)\right)$ (12)
+$\operatorname{tr}\left(B\left(X\right)Y\right)$ $=\operatorname{tr}\left(A\left(TXT^{\top}{}\right)A\left(Y\right)\right)$
 $=\operatorname{tr}\left(TXT^{\top}A\left(Y\right)\right)$
 $=\operatorname{tr}\left(XT^{\top}A\left(Y\right)T\right).$
-This shows that
-$B^{*}\left(Y\right)=T^{\top}A\left(Y\right)T.$
-It then follows that if
+This shows that $B^{*}\left(Y\right)=T^{\top}A\left(Y\right)T$ as required. To show the second claim observe that if
 $\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*},$
 then there exists a
-$\begin{bmatrix}Y_{1}&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}=\begin{bmatrix}\operatorname{diag}_{m}\left(Y_{1}\right)&\operatorname{diag}_{m}\left(Y_{2}\right)\\ \operatorname{diag}\left(Y_{2}^{\top}\right)&\operatorname{diag}_{m}\left(Y_{3}\right)\end{bmatrix}\succeq{}0$
+$\begin{bmatrix}Y_{1}&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}=\begin{bmatrix}\operatorname{diag}_{m}\left(Y_{1}\right)&\operatorname{diag}_{m}\left(Y_{2}\right)\\ \operatorname{diag}_{m}\left(Y_{2}^{\top}\right)&\operatorname{diag}_{m}\left(Y_{3}\right)\end{bmatrix}\succeq{}0$
 such that
-$\begin{bmatrix}0&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}0&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}T.$
+$\begin{bmatrix}0&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}0&Y_{1}\\ Y_{1}^{\top}&Y_{2}\end{bmatrix}T.$
 It is then straightforward to show that
-$\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}Y_{2}Y_{3}^{+}Y_{2}^{\top}&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}T,$
-which is in $\mathcal{P}^{*}$ as required. ∎
-We now prove 3.
-Proof.
-Observe that by 1($i$), the optimization in 3 is lower bounded by the relaxed optimization problem
-Minimize $\sum_{t=0}^{\infty}\operatorname{tr}\left(\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}{}&R_{2}\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\right)$
-over $\{U_{1}(t),U_{2}(t)\}_{t=0}^{\infty},$
-subject to $X(t+1)=\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix},$
-$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\in\mathcal{P}_{Q},X(0)\succeq{}0.$
-An analogous argument to that made in section 3.1 and 1($ii$) then shows that the set
-$\mathcal{C}_{\mu}=\left\{\Lambda:\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}\right\}$
-is non-empty, it has a minimum element.
-Therefore under (a) and (b) in the theorem statement and by 2, the relaxed optimization problem meets all the conditions of 1, from which it soon follows that the relaxed optimization has a finite value if and only if there exists a
-$\begin{bmatrix}\Lambda^{*}&0\\ 0&0\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
-such that
-$\Lambda^{*}=S+F^{\top}\Lambda^{*}F-\left(F^{\top}\Lambda^{*}G+R_{1}\right)\left(R_{2}+G^{\top}{}\Lambda^{*}G\right)^{+}\left(G^{\top}\Lambda^{*}F+R_{1}^{\top}{}\right).$
-Again by 1, the elements of the optimal input sequence $\overline{U}_{1}(t)$ and $\overline{U}_{2}(t)$ are the minimizers of the following optimization problem
-Minimize $\operatorname{tr}\left(\begin{bmatrix}0&R_{1}+F^{\top}\Lambda^{*}G\\ R_{1}^{\top}+G^{\top}\Lambda^{*}F&R_{2}+G^{\top}\Lambda^{*}G\end{bmatrix}\begin{bmatrix}0&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\right)$
-over $U_{1},U_{2},$
-subject to $\begin{bmatrix}X(t)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\in\mathcal{P}_{Q}.$
-The result will then follow if
-$\begin{bmatrix}X(t)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\succeq{}0,$
-since then all inputs to the relaxed problem are feasible also for the unrelaxed problem.
-It follows from an argument analogous to that in eq.
-12, and from the assumption that $X(0)\succeq{}0$, that we may replace the constraint in the above with
-$\begin{bmatrix}X(0)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\succeq{}0$ (13)
-without affecting the optimal value. Just as for the standard LQR problem it then follows from complementary slackness that there exists a $K$ such that
-$\begin{bmatrix}I\\ K^{\top}\end{bmatrix}\Lambda^{*}\begin{bmatrix}I&K\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
-and
-$\begin{bmatrix}I\\ K^{\top}\end{bmatrix}\Lambda^{*}\begin{bmatrix}I&K\end{bmatrix}\begin{bmatrix}X(0)&\overline{U}_{1}\\ \overline{U}_{1}^{\top}&\overline{U}_{2}\end{bmatrix}=0.$ (14)
-Applying this control results in $X(1)\succeq{}0$, and rerunning this argument for all $t$ then shows that eqs.
-13 and 14 hold for all $t$, not just $t=0$.
-Hence the optimal actions for the relaxed problem are feasible also for the unrelaxed problem, from which the result follows.
-∎
+$\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}Y_{2}Y_{2}^{+}Y_{1}^{\top}&Y_{1}\\ Y_{1}^{\top}&Y_{2}\end{bmatrix}T,$
+which is in $\mathcal{P}^{*}_{Q}$. Since $\mathcal{P}_{Q}^{*}$ is contained in the semi-definite cone, $M_{1}M_{2}^{+}M_{1}^{\top}$ is the minimum element of the set
+$\mathcal{C}_{\mu}=\left\{\Lambda:\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*}_{Q}\right\},$
+and so the formula for the infimum follows by 2. In order to verify the equation for the minimizer, first observe that from the above,
+$A\left(T\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}X\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}^{\top}{}T^{\top}\right)=$
+$\quad\begin{bmatrix}I&0\\ -Y_{2}^{+}Y_{1}^{\top}{}&0\end{bmatrix}A\left(T\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}T^{\top}\right)\begin{bmatrix}I&0\\ -Y_{2}^{+}Y_{1}^{\top}{}&0\end{bmatrix}^{\top}\!\!\!\!.$
+It then follows that
+$\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\in\mathcal{P}_{Q}\implies\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}X\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}^{\top}{}\!\!\!\!\!\!\in\mathcal{P}_{Q}.$
+Next observe that since $\left(I-M_{2}^{+}M_{2}\right)M_{1}^{\top}=0$ (else there would exist no $\Lambda$ such that eq. 12 holds),
+$\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}\!\!&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\!\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}\!X\!\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}^{\top}{}\!\!\!\!\!\!=0.$ (16)
+Finally for fixed $X$ note that the minimizer for eq. 13 equals the minimizer for
+$\inf\!\left\{\!\operatorname{tr}\!\left(\!\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}\!\!\!&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\!\!\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\!\right)\!:\!\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\!\in\mathcal{P}_{Q}\!\right\}\!.$
+From the definition of a dual cone,
+$\operatorname{tr}\!\left(\!\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\right)\geq{}0.$
+Therefore by eq. 16 the right hand side of eq. 14 is a minimizer as required. ∎